# Correspondence between two exact gravitational lens equations in a static and spherically symmetric spacetime

### v1

@@ Abstract
Virbhadra and Ellis have proposed an exact equation (referred to as VE equation) for the gravitational lens in a static and spherically symmetric spacetime [Phys. Rev. D 62,084003 (2000)], whereas an improved equation (referred to as BTOA equation) has been derived by Bozza [Phys. Rev. D 78, 103005 (2008)] and later by Takizawa, Ono and Asada [Phys. Rev. D, 102, 064060 (2020)]. VE and BTOA equations seem to be very different from each other. The present paper shows that there exists an unphysical branch in VE equation. Consequently, VE equation can be reduced by removing the unphysical branch. The reduced version of VE equation is found to be the same as BTOA equation when a suitable transformation is made between the deflection angles defined differently in the two formulations. An explicit expression of the transformation is found. We also argue possible numerical errors when the transformation between the deflection angles is ignored.
@@ I Introduction
As a pioneering work, Virbhadra and Ellis (VE) have proposed an exact equation for the gravitational lens in a static and spherically symmetric spacetime VE . Their geometrical approach is almost exact, though they assume that the deflection of light is caused on the lens plane. Their assumption is a good approximation when the observer, lens and source configuration is almost symmetric. Recently, VE equation has been used for detailed discussions of image distortions of Schwarzschild lensing Virbhadra2022 ; Virbhadra2024 .
Instead of assuming such a symmetrical configuration, Bozza has made use of the Ohanian relation for a more general configuration Ohanian to obtain an improved version of the exact lens equation Bozza2008 . Takizawa, Ono and Asada have later rediscovered this equation by using a state-of-art formulation GW ; Ishihara2016 ; Ishihara2017 based on the Gauss-Bonnet theorem Takizawa2020 . For its simplicity, the improved equation is referred to as BTOA equation. BTOA equation seems to be very different from VE equation. Is it true?
The main purpose of this paper is to examine whether VE equation is really different from BTOA one. First, we shall show that VE equation has a duality that it has not only a physical branch but also an unphysical one. Hence, VE equation can be reduced by removing the unphysical branch. Secondly, we shall argue that the reduced version of VE equation can be the same as BTOA equation when we make a suitable transformation between the deflection angles defined differently in the two formulations.
This paper is organized as follows. Section II discusses the unphysical branch in VE equation. Section III clarifies a relation between the reduced VE equation and BTOA one. Section IV argues possible numerical errors when the deflection angle transformation is ignored. Throughout this paper, we focus on a static and spherically symmetric spacetime.
II Two exact lens equations
II.1 VE and BTOA equations
Let us begin with briefly summarizing VE and BTOA equations. Both equations describe the mapping between the source angle $\beta$ and image angle $\theta$ in terms of angular diameter distances in rigorous manners (e.g. See Figure 1). $D_{L}$, $D_{S}$ and $D_{LS}$ denote the angular diameter distances between the observer and the lens plane, between the observer and the source plane, and between the lens and the source plane, respectively.
A crucial difference between the two equations is how to treat the deflection angles of light. See Figure 1. VE equation assumes a symmetric configuration in which the observer and source are equidistant from the lens, as usually assumed in the conventional lens equation with the thin lens approximation SEF .
@@ II.1 VE and BTOA equations
In order to avoid an almost symmetric assumption, Ohanian considered the tangent line at the source and that at the observer (See Figure 2) and the intersection point Q Ohanian . Note that the intersection point Q is generally off the lens plane. For the quadrilateral OLSQ, a relation among the four interior angles can be obtained. It is Ohanian relation Ohanian . Bozza rewrote it in terms of the angular diameter distances to arrive at an improved version of an exact lens equation as
$D_{S}\tan\beta=\frac{D_{L}\sin\theta-D_{LS}\sin(\alpha_{B}-\theta)}{\cos(\alpha_{B}-\theta)},$ (2)
where $\alpha_{B}$ is defined at the intersection point Q in Figure 2.
Figure 2: The deflection point Q in the BTOA equation is off the lens plane.
By using the Gauss-Bonnet method on the gravitational lens GW , another exact lens equation is derived as Takizawa2020
@@
where $\alpha_{TOA}$ is defined at the intersection point Q in Figure 2. This equation is equivalent to Eq. (2). See Reference Takizawa2020 for a proof of the equivalence. Hence, $\alpha_{B}=\alpha_{TOA}$.
@@ II.2 Duality in VE equation
where $A\equiv(D_{S})^{2}\tan^{2}\beta+4D_{L}D_{LS}$, $B\equiv D_{S}(D_{LS}-D_{L})\tan\beta$, and $A>0$. Eq. (7) tells ${\cal D}>0$ for any $\alpha_{VE}$.
@@ II.5 Reduced VE equation
III Correspondence between VE and BTOA equations
III.1 Transformation between the VE and BTOA deflection angles
For the same configuration of the lens, observer and source with given $\beta$ and $\theta$, the reduced VE and BTOA equations should equal to each other. For Eq. (11) to equal to Eq. (2), a relation between $\alpha_{VE}$ and $\alpha_{B}$ must be
@@ III.1 Transformation between the VE and BTOA deflection angles
where functions $E$ and $F$ of $\alpha_{B}$ are defined as
$E$ $\equiv D_{LS}\sin\alpha_{B}+D_{S}\tan\beta\cos\alpha_{B},$ (14)
$F$ $\equiv D_{L}+D_{LS}\cos\alpha_{B}-D_{S}\tan\beta\sin\alpha_{B}.$ (15)
Figure 7 shows a relation between $\alpha_{VE}$ and $\alpha_{B}$ in Eq. (13).
It is worthwhile to mention that BTOA equation as Eq. (2) can be simply rewritten in terms of $E$ and $F$ as
@@
Figure 7: $\tan\alpha_{VE}$ vs $\tan\alpha_{B}$. Top: $\beta=0$, and $D_{LS}/D_{L}=$ 1 (blue), 0.5 (orange), 0.3 (green) and 0.1 (red). Bottom: The equidistance case. $\beta=$ 0 (blue), 0.1 (orange), 0.2 (green) and 0.3 (red).
IV Numerical errors by ignoring the deflection angle transformation
@@
$\tan\alpha_{VE}-\tan\alpha_{B}=O\left(\frac{D_{L}}{D_{LS}}\right).$ (17)
Hence, $\alpha_{VE}$ is very close to $\alpha_{B}$. This means that the substitution of $\alpha_{B}$ into $\alpha_{VE}$ in Eq. (1) can be a good approximation even if the transformation is neglected.
@@
$\tan\alpha_{VE}-\tan\alpha_{B}=-\frac{1}{\tan\beta\cos\alpha_{B}}+O\left(\frac{D_{LS}}{D_{L}}\right).$ (18)
This implies that the difference between $\alpha_{VE}$ and $\alpha_{B}$ can be significantly large in the limit as $D_{LS}/D_{L}\to 0$, especially for nearly coaligned configurations $\beta\sim 0$. This case of $D_{L}\gg D_{LS}$ and $\beta\sim 0$ are corresponding to the observation of photons from galactic centers, notably a vicinity of a photon surface of the massive black hole candidate in Sgr A.
Figure 8: $\tan\theta_{VE}$ and $\tan\theta_{B}$. Here, the transformation of $\alpha_{VE}$ is ignored when we solve the reduced VE equation. The dashed line corresponds to $\tan\theta_{VE}=\tan\theta_{B}$ as a reference. $D_{LS}/D_{L}=10^{-1}$ (blue), $10^{-2}$ (red), and $10^{-3}$ (green), where $\beta=0.1$ is chosen.
How large numerical errors are caused when we ignore the transformation between the two deflection angles? For this purpose, we intend to equate $\alpha_{VE}$ and $\alpha_{B}$. We focus on the case of $|\tan\alpha_{VE}|\ll 1$ and $|\tan\alpha_{B}|\ll 1$. From Eqs. (1) and (2), then, we can obtain
$\cfrac{\tan\theta_{VE}-\tan\beta}{\cfrac{D_{LS}}{D_{S}}(1+\tan^{2}\beta)}=\cfrac{\tan\theta_{B}-\tan\beta}{\cfrac{D_{LS}}{D_{S}}+\tan^{2}\beta}+O(\tan^{2}\alpha_{VE},\tan^{2}\alpha_{B}),$ (19)
where $\theta_{VE}$ and $\theta_{B}$ denote a solution for Eqs. (1) and (2), respectively. We ignore the second order in $\tan\alpha_{B}$ below.
For $\tan\theta_{VE}$ to equal to $\tan\theta_{B}$, the two denominators in Eq. (19) must be the same as each other. However, it is not the case. See Figure 8 for numerical plots of Eq. (19).
The two denominators are very close to each other, if and only if $|\beta|\ll 1$ and $D_{L}\ll D_{LS}$. This condition means that the observer, lens and source are almost coaligned and the observer is close to the lens. This condition does not meet in most astronomical situations.
The exceptional case is the gravitational lens by the Sun. In many astronomical cases ($\beta\neq 0$ and $D_{L}\gg D_{LS}$), conversely speaking, the ignorance of the transformation between the deflection angles would result in a significant difference between $\tan\theta_{VE}$ and $\tan\theta_{B}$.
@@
$\tan\alpha_{VE}-\tan\alpha_{B}=\frac{D_{L}}{D_{LS}}\sin^{2}\beta\tan\alpha_{B}+O\left(\tan^{2}\alpha_{B}\right),$ (20)
which can be obtained from Eq. (13) in the approximation as $|\tan\alpha_{B}|\ll 1$. Eq. (20) is consistent with also Figure 7 in the domain of small $\tan\alpha_{B}$.
Figure 9: Numerical errors by ignoring the transformation between $\alpha_{VE}$ and $\alpha_{B}$ when the reduced VE equation is solved. The parameters correspond to the Sgr A Gillessen in the approximation of the Schwarzschild lens, where the angular radius of the black hole shadow PT is about 26 microarcsec., $D_{L}=8$ kpc, and $D_{LS}/D_{L}=1.2\times 10^{-10}$ EHT2022 . Here, $\delta\equiv|\theta_{VE}/\theta_{B}-1|$ is the relative difference due to neglecting the transformation.
@@ V Summary
We have shown that there exists an unphysical branch in VE equation. VE equation is reduced by removing the unphysical branch. The reduced VE equation has been found to be the same as BTOA equation when the transformation between the deflection angles is suitably done. We have argued possible numerical errors when the transformation is ignored. In particular, numerical errors can be significant in the use of the VE equation for the Sgr A.
@@ Acknowledgements.
We are grateful to Keita Takizawa and Marcus Werner for fruitful discussions. We thank Yuuiti Sendouda and Ryuichi Takahashi for useful conversations. This work was supported in part by Japan Science and Technology Agency (JST) SPRING, Grant Number, JPMJSP2152 (R.K.) and in part by Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research, No. 20K03963 and 24K07009(H.A.).

### v2

# Correspondence between two gravitational lens equations in a static and spherically symmetric spacetime

@@ Abstract
Virbhadra and Ellis have proposed a very accurate equation (referred to as VE equation) for the gravitational lens in a static and spherically symmetric spacetime [Phys. Rev. D 62,084003 (2000)], whereas an improved equation (referred to as OB equation) has been derived by Bozza [Phys. Rev. D 78, 103005 (2008)] based on a relation found by Ohanian [Am. J. Phys. 55, 428 (1987)]. OB equation was rediscovered later by Takizawa, Ono and Asada [Phys. Rev. D, 102, 064060 (2020)]. VE and OB equations seem to be very different from each other. The present paper shows that there exists an unphysical branch in VE equation. Consequently, VE equation can be improved by removing the unphysical branch. The improved version of VE equation is found to be the same as OB equation when a suitable transformation is made between the deflection angles defined differently in the two formulations. An explicit expression of the transformation is found. We also argue possible numerical differences when the transformation between the deflection angles is ignored.
@@ I Introduction
As a pioneering work, Virbhadra and Ellis (VE) have proposed a very accurate equation for the gravitational lens in a static and spherically symmetric spacetime VE . Their geometrical approach is almost exact Perlick , though they assume that the deflection of light is caused on the lens plane. Their assumption is a good approximation when the observer, lens and source configuration is almost aligned. Recently, VE equation has been used for detailed discussions of image distortions of Schwarzschild lensing Virbhadra2022 ; Virbhadra2024 .
Instead of assuming such a symmetrical configuration, Bozza has made use of the Ohanian relation for a more general configuration Ohanian to obtain an improved version of the accurate lens equation Bozza2008 . For its simplicity, the improved equation is referred to as Ohanian-Bozza (OB) equation. Takizawa, Ono and Asada have later rediscovered this equation in an apparently different form by using a state-of-art formulation GW ; Ishihara2016 ; Ishihara2017 based on the Gauss-Bonnet theorem Takizawa2020 . OB equation seems to be very different from VE equation. Is it true?
The main purpose of this paper is to examine whether VE equation is really different from OB one. First, we shall show that VE equation has a duality that it has not only a physical branch but also an unphysical one. Hence, VE equation can be improved by removing the unphysical branch. Secondly, we shall argue that the improved version of VE equation can be the same as OB equation when we make a suitable transformation between the deflection angles defined differently in the two formulations.
This paper is organized as follows. Section II discusses the unphysical branch in VE equation. Section III clarifies a relation between the improved VE equation and OB one. Section IV argues possible numerical differences when the deflection angle transformation is ignored. Throughout this paper, we focus on a static and spherically symmetric spacetime.
II Two lens equations
II.1 VE and OB equations
Let us begin with briefly summarizing VE and OB equations. Both equations describe the mapping between the source angle $\beta$ and image angle $\theta$ in terms of angular diameter distances in rigorous manners (e.g. See Figure 1). $D_{L}$, $D_{S}$ and $D_{LS}$ denote the angular diameter distances between the observer and the lens plane, between the observer and the source plane, and between the lens and the source plane, respectively.
A crucial difference between the two equations is how to treat the deflection angles of light. See Figure 1. VE equation assumes an almost aligned configuration in which the observer and source are equidistant from the lens, as usually assumed in the conventional lens equation with the thin lens approximation SEF .
@@ II.1 VE and OB equations
In order to avoid an almost aligned assumption, Ohanian considered the tangent line at the source and that at the observer (See Figure 2) and the intersection point Q Ohanian . Note that the intersection point Q is generally off the lens plane. For the quadrilateral OLSQ (See Figure 2), a relation among the four interior angles can be obtained. It is Ohanian relation Ohanian . Bozza rewrote it in terms of the angular diameter distances to arrive at an improved version of an accurate lens equation as
$D_{S}\tan\beta=\frac{D_{L}\sin\theta-D_{LS}\sin(\alpha_{OB}-\theta)}{\cos(\alpha_{OB}-\theta)},$ (2)
where $\alpha_{OB}$ is defined at the intersection point Q in Figure 2.
Figure 2: The deflection point Q in the OB equation is off the lens plane.
By using the Gauss-Bonnet method on the gravitational lens GW , another accurate lens equation is derived as Takizawa2020
@@
where $\alpha_{TOA}$ is defined at the intersection point Q in Figure 2. This equation is equivalent to Eq. (2). See Reference Takizawa2020 for a proof of the equivalence. Hence, $\alpha_{OB}=\alpha_{TOA}$.
@@ II.2 Duality in VE equation
where $A\equiv(D_{S})^{2}\tan^{2}\beta+4D_{L}D_{LS}$, $B\equiv D_{S}(D_{LS}-D_{L})\tan\beta$, and $A>0$. Eq. (7) gives ${\cal D}>0$ for any $\alpha_{VE}$.
@@ II.5 Improved VE equation
III Correspondence between VE and OB equations
III.1 Transformation between the VE and OB deflection angles
For the same configuration of the lens, observer and source with given $\beta$ and $\theta$, the improved VE and OB equations should equal to each other. For Eq. (11) to equal to Eq. (2), a relation between $\alpha_{VE}$ and $\alpha_{OB}$ must be
@@ III.1 Transformation between the VE and OB deflection angles
where functions $E$ and $F$ of $\alpha_{OB}$ are defined as
$E$ $\equiv D_{LS}\sin\alpha_{OB}+D_{S}\tan\beta\cos\alpha_{OB},$ (14)
$F$ $\equiv D_{L}+D_{LS}\cos\alpha_{OB}-D_{S}\tan\beta\sin\alpha_{OB}.$ (15)
Figure 7 shows a relation between $\alpha_{VE}$ and $\alpha_{OB}$ in Eq. (13).
It is worthwhile to mention that OB equation as Eq. (2) can be simply rewritten in terms of $E$ and $F$ as
@@
Figure 7: $\tan\alpha_{VE}$ vs $\tan\alpha_{OB}$. Top: $\beta=0$, and $D_{LS}/D_{L}=$ 1 (blue), 0.5 (orange), 0.3 (green) and 0.1 (red). Bottom: The equidistance case. $\beta=$ 0 (blue), 0.1 (orange), 0.2 (green) and 0.3 (red).
IV Numerical differences by ignoring the deflection angle transformation
@@
$\tan\alpha_{VE}-\tan\alpha_{OB}=O\left(\frac{D_{L}}{D_{LS}}\right).$ (17)
Hence, $\alpha_{VE}$ is very close to $\alpha_{OB}$. This means that the substitution of $\alpha_{OB}$ into $\alpha_{VE}$ in Eq. (1) can be a good approximation even if the transformation is neglected.
@@
$\tan\alpha_{VE}-\tan\alpha_{OB}=-\frac{1}{\tan\beta\cos\alpha_{OB}}+O\left(\frac{D_{LS}}{D_{L}}\right).$ (18)
This implies that the difference between $\alpha_{VE}$ and $\alpha_{OB}$ can be significantly large in the limit as $D_{LS}/D_{L}\to 0$, especially for nearly coaligned configurations $\beta\sim 0$. This case of $D_{L}\gg D_{LS}$ and $\beta\sim 0$ are corresponding to the observation of photons from centers of galaxies, notably a vicinity of a photon surface of the massive black hole candidate in Sgr A.
Figure 8: $\tan\theta_{VE}$ and $\tan\theta_{OB}$. Here, the transformation of $\alpha_{VE}$ is ignored when we solve the improved VE equation. The dashed line corresponds to $\tan\theta_{VE}=\tan\theta_{OB}$ as a reference. $D_{LS}/D_{L}=10^{-1}$ (blue), $10^{-2}$ (red), and $10^{-3}$ (green), where $\beta=0.1$ is chosen.
How large are numerical differences when we ignore the transformation between the two deflection angles? For this purpose, we intend to equate $\alpha_{VE}$ and $\alpha_{OB}$. We focus on the case of $|\tan\alpha_{VE}|\ll 1$ and $|\tan\alpha_{OB}|\ll 1$. From Eqs. (1) and (2), then, we can obtain
$\cfrac{\tan\theta_{VE}-\tan\beta}{\cfrac{D_{LS}}{D_{S}}(1+\tan^{2}\beta)}=\cfrac{\tan\theta_{OB}-\tan\beta}{\cfrac{D_{LS}}{D_{S}}+\tan^{2}\beta}+O(\tan^{2}\alpha_{VE},\tan^{2}\alpha_{OB}),$ (19)
where $\theta_{VE}$ and $\theta_{OB}$ denote a solution for Eqs. (1) and (2), respectively. We ignore the second order in $\tan\alpha_{OB}$ below.
For $\tan\theta_{VE}$ to equal to $\tan\theta_{OB}$, the two denominators in Eq. (19) must be the same as each other. However, it is not the case. See Figure 8 for numerical plots of Eq. (19).
The two denominators are very close to each other, if and only if $|\beta|\ll 1$ and $D_{L}\ll D_{LS}$. This condition means that the observer, lens and source are almost coaligned and the observer is close to the lens. This condition is not satisfied for most of astronomical situations.
The exceptional case is the gravitational lens by the Sun. In many astronomical cases ($\beta\neq 0$ and $D_{L}\gg D_{LS}$), conversely speaking, the ignorance of the transformation between the deflection angles would result in a significant difference between $\tan\theta_{VE}$ and $\tan\theta_{OB}$.
@@
$\tan\alpha_{VE}-\tan\alpha_{OB}=\frac{D_{L}}{D_{LS}}\sin^{2}\beta\tan\alpha_{OB}+O\left(\tan^{2}\alpha_{OB}\right),$ (20)
which can be obtained from Eq. (13) in the approximation as $|\tan\alpha_{OB}|\ll 1$. Eq. (20) is consistent with also Figure 7 in the domain of small $\tan\alpha_{OB}$.
Figure 9: Numerical differences by ignoring the transformation between $\alpha_{VE}$ and $\alpha_{OB}$ when the improved VE equation is solved. The parameters correspond to the Sgr A Gillessen in the approximation of the Schwarzschild lens, where the angular radius of the black hole shadow PT is about 26 microarcsec., $D_{L}=8$ kpc, and $D_{LS}/D_{L}=1.2\times 10^{-10}$ EHT2022 . Here, $\delta\equiv|\theta_{VE}/\theta_{OB}-1|$ is the relative difference due to neglecting the transformation.
@@ V Summary
We have shown that there exists an unphysical branch in VE equation. VE equation is improved by removing the unphysical branch. The improved VE equation has been found to be the same as OB equation when the transformation between the deflection angles is suitably done. We have argued possible numerical differences when the transformation is ignored. In particular, numerical differences can be significant in the use of the VE equation for the Sgr A.
@@ Acknowledgements.
We thank Shwetketu Virbhadra for comments on the earlier version of our manuscript. We wish to thank an anonymous referee for helpful comments. We are grateful to Keita Takizawa and Marcus Werner for fruitful discussions. We thank Yuuiti Sendouda and Ryuichi Takahashi for useful conversations. This work was supported in part by Japan Science and Technology Agency (JST) SPRING, Grant Number, JPMJSP2152 (R.K.) and in part by Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research, No. 20K03963 and 24K07009(H.A.).